# references.bib

### v1

@@ 3 Discussion
We derived a reparametrisation of Totimorphic lattices in two and three dimensions that enables the use of constraint-free gradient descent to optimise the lattice topology. Different from previously proposed mechanically reprogrammable systems [lee2022mechanical, liu2022triclinic], the effective physical properties of Totimorphic lattices can be changed continuously.
This was successfully illustrated using the introduced framework for two space-related applications: continuous control of the mechanical properties of a metamaterial as well as continuous control of the optical properties of a conceptual telescope mirror composed of Totimorphic unit cells.
Instead of reparametrising the lattice, a variety of other techniques can be applied as well [platt1987constrained]. As an example, for gradient-based optimisation the penalty method can be used where the Totimorphic condition is added as a penalty term in the cost function to be reinforced. However, it comes with the downside that the constraints are not perfectly satisfied during optimisation (hence we do not obtain trajectories), and in practice it can be non-trivial to avoid becoming stuck in local minima when trying to balance the penalty and the objective-dependent terms in the cost function. Another alternative is the method of Lagrange multipliers, but it does not work with gradient descent, requiring more involved solvers capable of finding saddle points of the cost function. An extension of the Lagrange multiplier method, called basic differential multiplier method [platt1987constrained], which is compatible with gradient descent can be used instead, although – as in the case of the penalty method – constraints are not satisfied throughout the whole optimisation process. Apart from methods using gradient descent, more advanced approaches that divide the problem into simpler sub-problems can be used, such as sequential quadratic programming [schittkowski1986nlpql]. In this work, we stick to the gradient-based approach, as it combines ease of formulation with a straightforward and flexible algorithmic implementation, which is scaleable to large lattices and more involved cost functions, e.g., including non-linear functions of the physical lattice coordinates such as deep neural networks. Most importantly, it defines proper dynamics driven by an objective-dependent cost function to reconfigure the lattice – something most other methods are lacking.
For the model in two dimensions, we were able to formulate simple rules for detecting when lattice elements collide during reconfiguration. In contrast, we did not employ any collision detection methods for the three-dimensional model, restricting our experiments to cases where strong folding of the lattice surface does not occur. In future work, approaches used to simulate textiles [baraff2023large] could be employed to detect self-collisions in Totimorphic structures. Apart from simply detecting whether the structure collides with itself, it is even more important to deal with self-collisions and invalid lattice configurations [^7] I.e., values of the generalised coordinates that yield no valid solution for the physical lattice coordinates. during the optimisation process. A default method for dealing with collisions in, e.g., simulations of swarming robots [izzo2007autonomous] or textile simulations [baraff2023large] is to add strong repulsive forces that only act between elements at very low distances. Adopting such an approach for Totimorphic structures is certainly feasible, but might be problematic due to the rigid and discontinuous nature of lattices. Furthermore, this method only helps in dealing with self-collisions and increases the computational cost of simulations substantially. Alternatively, one could truncate parameter updates during optimisation such that general invalid lattice configurations are avoided [^8] As proposed in Matthew Fisher’s blog post on cloth self-collisions, https://graphics.stanford.edu/~mdfisher/cloth.html, accessed 2024. – although the optimisation process can get stuck trying to pass through invalid configurations, requiring either stochasticity to escape such local minima of the cost function or intermixing with optimisers that allow discontinuous changes of the lattice configuration. In the presented experiments, we only checked whether the lattice is still valid after each update and stopped in case the lattice broke.
Even though we only derived the model for a specific tiling here, the approach can be generalized to arbitrary Totimorphic structures. It further enables a theoretical study of Totimorphic designs. For example, it is an open question whether the Totimorphic configuration space is fully connected or consists of topologically disconnected sub-spaces; or put differently: it is not clear whether there always exists a continuous trajectory connecting two valid Totimorphic configurations. This is particularly important when utilising optimisation methods that – different to gradient descent – facilitate a global exploration of the configuration space, such as evolutionary strategies, to avoid selecting solutions that are not reachable from the initial lattice configuration.
In the first application, we demonstrated how a Totimorphic lattice can be reconfigured to continuously change it’s Poisson’s ratio from negative to positive values and vice versa. This was illustrated for two-dimensional truss lattices only. For three dimensional metamaterials, the same approach can be used by exchanging the two-dimensional Totimorphic model and direct stiffness method with their three-dimensional equivalents. However, appropriate volume-filling Totimorphic tilings will have to be designed to achieve similar results than in the two-dimensional case.
@@
In the second application, we demonstrated how a Totimorphic lattice can be used as the backbone for autonomously deployable and reconfigurable large-scale space infrastructure, specifically the primary mirror element of a telescope. A major criterion for the operability of such a mirror is how accurately it focuses the reflected light, as this directly limits the attainable resolution when imaging objects. For objects at far larger distances, e.g., with angular sizes similar to Jupiter, we can already see that the magnification of our design is not sufficient to produce sharp images – although rough details are still recognizable (see Fig. S3). This can be improved by further increasing the focal length. An approach that is often used in modern telescopes is to guide the light through a hierarchy of mirrors, e.g., by adding a secondary or even tertiary mirror before projecting the light onto a sensor. In our case, those mirrors would be Totimorphic as well. Due to the differentiability of the whole process, our framework can be used to (re)configure all mirrors simultaneously. This way, subsequent mirrors can even compensate for distortions introduced by previous (or later) mirrors. In addition, the focusing of a mirror based on a Totimorphic lattice can be improved by reducing the size of unit cells. The effect of miniaturisation on focal length and the physical properties of the Totimorphic cells must be quantified to mature this concept.
@@
Developing a practical prototype of a Totimorphic structure will require overcoming many pressing technical challenges, as the underlying assumptions and abstractions used in the model become physically unrealistic. For instance, it is assumed that the beams and levers are perfectly rigid. In reality, deformations in the structure from external or internal forces will produce bending in the beams, destabilising the internal spring forces. Similarly, the zero-length property of the springs will likely not hold for a real structure. While it is possible to create springs that approximate the property for a range of lengths, at small or large extensions the zero-length property will likely not persist. Moreover, when in the zero-stiffness state, the Totimorphic structure should require no work to move between configurations, but practically there will be friction and jamming in the joints of the structure, which requires work from actuators to overcome. Independent of the application, minimising the number of actuators required to operate a spacecraft is quite critical, as each actuator represents a mode of failure that might compromise the spacecraft. Vacuum-exposed mechanisms and actuators are also at risk of cold welding and are difficult to lubricate. Since the number of actuators required to control a Totimorphic lattice scales linearly with the number of unit cells, it will be critical to investigate whether it is possible to reduce the number of actuators and optimise the operations required to morph between given configurations. The specifics of transforming 1D springs, Euler beams and frictionless joints into a manufacturable model is beyond the scope of this paper, but it will be necessary to make decisions that balance the requirements with suitability for space on various aspects like material choices, beam/lever aspect ratios, joint types, and actuation methods.
@@
Despite the challenges, Totimorphic lattices have a number of advantages for use in space environment. Deployable, reconfigurable, low-density structures are highly desirable for the strict mass and energy requirements of most space missions. We suggest that Totimorphic structures potentially fill a functional gap between Origami and inflatable structures, being able to carry tensile and compressive loads while also being flexible and capable of supporting surfaces without requiring pressurisation. Totimorphic structures may be compatible with further applications – a variety of space infrastructure use-cases such as space antennas and solar sails.
@@
To conclude, Totimorphic structures represent an intriguing class of mechanically reprogrammable systems, featuring continuously adjustable effective physical properties through simple mechanical actuations. Especially since our framework is light-weight and flexible regarding the cost function, it can be implemented on edge devices to allow autonomous and lattice-aware control of the actuators, e.g., using a deep neural network to estimate effective properties of the Totimorphic structure from the state of its actuators (the generalized coordinates). Thus, we are confident that the introduced concept can be scaled up to real-world, large-scale prototypes capable of autonomously and continuously adapting their lattice configuration to achieve a specified objective.
@@ 4 Methods
In the following, we describe the mathematical models referenced in the main text. These are just one particular realisation of an analytical Totimorphic model, and the introduced framework can be extended to other Totimorphic structures in a straightforward way. Implementations of the models in Python are available on Gitlab [^9] https://gitlab.com/EuropeanSpaceAgency/lattymorph.
@@ 4.1 Notation
$\boldsymbol{\epsilon}^{\alpha,\beta}=\begin{pmatrix}\cos{\alpha}\cos{\beta}\\ \sin{\beta}\\ -\sin{\alpha}\cos{\beta}\end{pmatrix}=R_{y}(\alpha)R_{z}(\beta)\mathbb{1}_{x}$ (13)
@@
$\mathbb{1}_{x}=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$ (14)
@@ 4.2 Totimorphic lattices in two dimensions
4.2.1 Parametrisation of elementary cells
In the following, we drop the index $ij$ (cell at row $i$ and column $j$) for clarity. However, all vectors are always given per elementary cell (see main text). An elementary cell is constructed from two unit cells which are connected at their levers. An elementary cell in the Totimorphic lattice is parameterized by $(\boldsymbol{o},\varphi,\theta,\varphi^{\prime},\theta^{\prime})$, where $\boldsymbol{o}$ is the location of point $A$, $\varphi$ is the angle with the x-axis of the lower beam, and $\theta$ is the angle of the lower lever with respect to the lower beam. Primed variables represent the same quantities, but for the upper half of the elementary cell (i.e., the upper beam and lever).
@@ 4.2.1 Parametrisation of elementary cells
$\boldsymbol{r}^{\mathrm{A}}$ $=\boldsymbol{o}\,,$ (17a)
$\boldsymbol{r}^{\mathrm{B}}$ $=\boldsymbol{r}^{\mathrm{A}}+l_{\text{b}}\cdot\boldsymbol{e}^{\varphi}\,,$ (17b)
$\boldsymbol{r}^{\mathrm{P}}$ $=\boldsymbol{r}^{\mathrm{A}}+\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi}\,,$ (17c)
$\boldsymbol{r}^{\mathrm{C}}$ $=\boldsymbol{r}^{\mathrm{P}}+\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi+\theta}\,.$ (17d)
@@
$\boldsymbol{r}^{\mathrm{P^{\prime}}}=\boldsymbol{r}^{\mathrm{C}}+\frac{l_{\text{b}}}{2}\boldsymbol{e}^{\gamma^{\prime}}$ $\ \ \ \ \text{with}\ \ \ \ \gamma^{\prime}=\pi-\theta^{\prime}+\varphi^{\prime}\,,$ (18a)
$\boldsymbol{r}^{\mathrm{A^{\prime}}}$ $=\boldsymbol{r}^{\mathrm{P^{\prime}}}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi^{\prime}}\,,$ (18b)
$\boldsymbol{r}^{\mathrm{B^{\prime}}}=\boldsymbol{r}^{\mathrm{P^{\prime}}}+$ $\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi^{\prime}}=2\cdot\boldsymbol{r}^{\mathrm{P^{\prime}}}-\boldsymbol{r}^{\mathrm{A^{\prime}}}\,.$ (18c)
@@
$\boldsymbol{e}^{\varphi^{\prime}}$ $=\frac{1}{l_{\text{b}}}\cdot\left(\boldsymbol{r}^{\mathrm{B^{\prime}}}-\boldsymbol{r}^{\mathrm{A^{\prime}}}\right)\,,$ (19a)
$\varphi^{\prime}$ $=\text{sign}\left(e^{\varphi^{\prime}}_{y}\right)\cdot\text{arccos}\left(e^{\varphi^{\prime}}_{x}\right)\,,$ (19b)
@@
$\boldsymbol{e}^{\gamma^{\prime}}$ $=\frac{2}{l_{\text{b}}}\cdot\left(\boldsymbol{r}^{\mathrm{P^{\prime}}}-\boldsymbol{r}^{\mathrm{C}}\right)\,,$ (20a)
$\gamma^{\prime}$ $=\text{sign}\left(e^{\gamma^{\prime}}_{y}\right)\cdot\text{arccos}\left(e^{\gamma^{\prime}}_{x}\right)\,,$ (20b)
$\theta^{\prime}$ $=\pi-\gamma^{\prime}+\varphi^{\prime}\,,$ (20c)
with $\varphi^{\prime}$ given by Eq. 19b.
@@
$\alpha^{\prime}_{\mathrm{L}}{}$ $=\frac{\theta^{\prime}}{2}-\varphi^{\prime}\,,$ (21a)
$\alpha^{\prime}_{\mathrm{R}}{}$ $=\frac{\pi}{2}-\alpha^{\prime}_{\mathrm{L}}{}\,,$ (21b)
$\alpha_{\mathrm{L}}{}$ $=\frac{\theta}{2}+\varphi\,,$ (21c)
$\alpha_{\mathrm{R}}{}$ $=\frac{\pi}{2}-\alpha_{\mathrm{L}}{}\,.$ (21d)
@@
4.2.2 Connecting elementary cells

### v2

@@ 3 Discussion
We introduced a differentiable parametrisation of Totimorphic lattices in two and three dimensions that enables constraint-free, gradient-based reconfiguration. Unlike most previously proposed reprogrammable systems [12, 13], Totimorphic lattices can vary their effective properties continuously through geometry alone. Furthermore, by absorbing constraints into generalised coordinates, smooth trajectories can be generated using gradient descent instead of only final states. The method is simple to implement, compatible with modern automatic differentiation frameworks, and scalable to complex objectives, including cost functions that contain differentiable surrogate models for property prediction such as deep neural networks.
@@
In our first application, we showed that a Totimorphic lattice can be reconfigured to continuously shift its Poisson’s ratio between negative and positive values. This was illustrated for two-dimensional truss lattices. Extending this principle to three-dimensional materials would require designing volume-filling Totimorphic tilings and employing 3D stiffness methods, but the underlying approach remains the same. In our second application, we explored Totimorphic lattices as the backbone for autonomously deployable and reconfigurable large-scale space infrastructure, focusing on a primary telescope mirror. A key criterion for the operability of such a mirror is how accurately it focuses the reflected light, as this directly limits the attainable resolution when imaging objects. For objects at far larger distances, e.g., with angular sizes similar to Jupiter, we can already see that the magnification of our design does not yet achieve sharp imaging – although rough details are still recognisable (see Fig. S3 in the Supplemental Information). This can be improved by further increasing the focal length. An approach that is often used in modern telescopes is to guide the light through a hierarchy of mirrors, e.g., by adding a secondary or even tertiary mirror before projecting the light onto a sensor. In our case, those mirrors would be Totimorphic as well. Due to the differentiability of the whole process, our framework can be used to (re)configure all mirrors simultaneously. This way, subsequent mirrors can even compensate for distortions introduced by previous (or later) mirrors. In addition, the focusing of a mirror based on a Totimorphic lattice can be improved by reducing the size of unit cells. The effect of miniaturisation on focal length and the physical properties of the Totimorphic cells must be quantified to mature this concept.
Our current models also highlight methodological challenges. For the model in two dimensions, we formulated simple rules for detecting when lattice elements collide during reconfiguration. In contrast, we did not employ any collision detection methods for the three-dimensional model, restricting our experiments to cases where strong folding of the lattice surface does not occur. In future work, approaches used to simulate textiles [2] could be employed to detect self-collisions in Totimorphic structures. Apart from simply detecting whether the structure collides with itself, it is even more important to deal with self-collisions and invalid lattice configurations during the optimisation process. A default method for dealing with collisions in, e.g., simulations of swarming robots [10] or textile simulations [2] is to add strong repulsive forces that only act between elements at very low distances – also known as barrier functions. Adopting such an approach for Totimorphic structures is certainly feasible, but might be problematic due to the rigid and discontinuous nature of lattices. Furthermore, this method only helps in dealing with self-collisions and increases the computational cost of simulations substantially. Alternatively, one could truncate parameter updates during optimisation such that general invalid lattice configurations are avoided (as proposed in [6]) – although the optimisation process can get stuck trying to pass through invalid configurations, requiring either stochasticity to escape such local minima of the cost function or intermixing with optimisers that allow discontinuous changes of the lattice configuration. In the presented experiments, we only checked whether the lattice is still valid after each update and stopped in case the lattice broke, which mostly occurred when attempting to reach too extreme mechanical properties.
Model-wise, instead of reparametrising the lattice, a variety of other techniques can be applied as well [19]. As an example, for gradient-based optimisation the penalty method can be used where the beam and lever length and angle constraints are added as a penalty terms in the cost function to be reinforced. However, it comes with the downside that the constraints are not perfectly satisfied during optimisation (hence we do not obtain trajectories), and in practice it can be non-trivial to avoid becoming stuck in local minima when trying to balance the penalty and the objective-dependent terms in the cost function. Another alternative is the method of Lagrange multipliers, but it does not work with gradient descent, requiring more involved solvers capable of finding saddle points of the cost function. An extension of the Lagrange multiplier method, called basic differential multiplier method [19], which is compatible with gradient descent can be used instead, although – as in the case of the penalty method – constraints are not satisfied throughout the whole optimisation process. Apart from methods using gradient descent, more advanced approaches that divide the problem into simpler sub-problems can be used, such as sequential quadratic programming [23], although these are also not guaranteed to produce continuous trajectories.
@@
On the theory side, it is an open question whether there exists always a continuous trajectory connecting two valid Totimorphic configurations. This is particularly important when utilising optimisation methods that – different to gradient descent – facilitate a global exploration of the configuration space, such as evolutionary strategies, to avoid selecting solutions that are not reachable from the initial lattice configuration. On the practical side, developing an actual prototype of a Totimorphic structure will require assessing the abstractions assumed in our model. For instance, in simulations we assume that the beams and levers are perfectly rigid. In reality, deformations in the structure from external or internal forces will produce bending in the beams, destabilising the internal spring forces. Similarly, the zero-length property of the springs will likely not hold for a real structure. While it is possible to create springs that approximate the property for a range of lengths, at small or large extensions the zero-length property will likely not persist. Moreover, when in the zero-stiffness state, the Totimorphic structure should require no work to move between configurations, but practically there will be friction and jamming in the joints of the structure, which requires work from actuators to overcome. Independent of the application, minimising the number of actuators required to operate a spacecraft is quite critical, as each actuator represents a mode of failure that might compromise the spacecraft. Vacuum-exposed mechanisms and actuators are also at risk of cold welding and are difficult to lubricate. Since the number of actuators required to control a Totimorphic lattice scales linearly with the number of unit cells, it will be critical to investigate whether it is possible to reduce the number of actuators and optimise the operations required to morph between given configurations.The specifics of transforming 1D springs, Euler beams and frictionless joints into a manufacturable model is beyond the scope of this paper, but it will be necessary to make decisions that balance the requirements with suitability for space on various aspects like material choices, beam/lever aspect ratios, joint types, and actuation methods. In future studies, the presented model should be extended to include such effects to reduce the gap between theory and practice.
@@
Despite these challenges, Totimorphic lattices have a number of advantages for use in the space environment. Deployable, reconfigurable, low-density structures are highly desirable for the strict mass and energy requirements of most space missions. We suggest that Totimorphic structures potentially fill a functional gap between non-rigid deployables like fabrics and inflatables, and rigid materials like origami structures, being able to carry tensile and compressive loads while also being flexible and capable of supporting surfaces without requiring pressurisation. Unlike either inflatable or origami structures, Totimorphic lattices are highly reconfigurable and could fulfil many different functions during a single deployment. In particular, Totimorphic structures may be compatible with a variety of additional space infrastructure use-cases such as space antennas and solar sails.
@@
To conclude, Totimorphic structures represent an intriguing class of mechanically reprogrammable systems, featuring continuously adjustable effective physical properties through simple mechanical and reversible actuations. Especially since our framework is light-weight and flexible regarding the cost function, it can be implemented on edge devices to allow autonomous and lattice-aware control of the actuators, e.g., using a deep neural network to estimate effective properties of the Totimorphic structure from the state of its actuators (the generalised coordinates). Thus, we are confident that the introduced concept can be scaled up to real-world, large-scale prototypes capable of autonomously and continuously adapting their lattice configuration to achieve a specified objective.
@@ 4 Methods
In the following, we describe the mathematical models referenced in the main text, including mathematical proofs. These are just one particular realisation of an analytical Totimorphic model, and the introduced framework can be extended to other Totimorphic structures in a straightforward way. Implementations of the models in Python are available on Github [8]. Simulation details are provided at the end of this section.
@@ 4.1 Notation
$\boldsymbol{\epsilon}^{\alpha,\beta}=\begin{pmatrix}\cos{\alpha}\cos{\beta}\\ \sin{\beta}\\ -\sin{\alpha}\cos{\beta}\end{pmatrix}=R_{y}(\alpha)R_{z}(\beta)\boldsymbol{1}_{x}$ (13)
@@
$\boldsymbol{1}_{x}=\begin{pmatrix}1\\ 0\\ 0\end{pmatrix}$ (14)
@@ 4.2 Totimorphic lattices in two dimensions
4.2.1 Parametrisation of elementary cells.
In the following, we drop the index $ij$ (cell at row $i$ and column $j$) for clarity. However, all vectors are always given per elementary cell (see main text). An elementary cell is constructed from two unit cells which are connected at their levers. An elementary cell in the Totimorphic lattice is parameterised by $(\boldsymbol{o},\varphi,\theta,\varphi^{\prime},\theta^{\prime})$, where $\boldsymbol{o}$ is the location of point $A$, $\varphi$ is the angle with the x-axis of the lower beam, and $\theta$ is the angle of the lower lever with respect to the lower beam. Primed variables represent the same quantities, but for the upper half of the elementary cell (i.e., the upper beam and lever).
@@
$\boldsymbol{r}^{\mathrm{A}}$ $=\boldsymbol{o}\,$, (17)
$\boldsymbol{r}^{\mathrm{B}}$ $=\boldsymbol{r}^{\mathrm{A}}+l_{\text{b}}\cdot\boldsymbol{e}^{\varphi}\,$, (18)
$\boldsymbol{r}^{\mathrm{P}}$ $=\boldsymbol{r}^{\mathrm{A}}+\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi}\,$, (19)
$\boldsymbol{r}^{\mathrm{C}}$ $=\boldsymbol{r}^{\mathrm{P}}+\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi+\theta}\,.$ (20)
@@
$\boldsymbol{r}^{\mathrm{P^{\prime}}}=\boldsymbol{r}^{\mathrm{C}}+\frac{l_{\text{b}}}{2}\boldsymbol{e}^{\gamma^{\prime}}$ $\ \ \ \ \text{with}\ \ \ \ \gamma^{\prime}=\pi-\theta^{\prime}+\varphi^{\prime}\,$, (21)
$\boldsymbol{r}^{\mathrm{A^{\prime}}}$ $=\boldsymbol{r}^{\mathrm{P^{\prime}}}-\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi^{\prime}}\,$, (22)
$\boldsymbol{r}^{\mathrm{B^{\prime}}}=\boldsymbol{r}^{\mathrm{P^{\prime}}}+$ $\frac{l_{\text{b}}}{2}\cdot\boldsymbol{e}^{\varphi^{\prime}}=2\cdot\boldsymbol{r}^{\mathrm{P^{\prime}}}-\boldsymbol{r}^{\mathrm{A^{\prime}}}\,.$ (23)
@@
$\boldsymbol{e}^{\varphi^{\prime}}$ $=\frac{1}{l_{\text{b}}}\cdot\left(\boldsymbol{r}^{\mathrm{B^{\prime}}}-\boldsymbol{r}^{\mathrm{A^{\prime}}}\right)\,$, (24)
$\varphi^{\prime}$ $=\text{sign}\left(e^{\varphi^{\prime}}_{y}\right)\cdot\text{arccos}\left(e^{\varphi^{\prime}}_{x}\right)\,$, (25)
@@
$\boldsymbol{e}^{\gamma^{\prime}}$ $=\frac{2}{l_{\text{b}}}\cdot\left(\boldsymbol{r}^{\mathrm{P^{\prime}}}-\boldsymbol{r}^{\mathrm{C}}\right)\,$, (26)
$\gamma^{\prime}$ $=\text{sign}\left(e^{\gamma^{\prime}}_{y}\right)\cdot\text{arccos}\left(e^{\gamma^{\prime}}_{x}\right)\,$, (27)
$\theta^{\prime}$ $=\pi-\gamma^{\prime}+\varphi^{\prime}\,$, (28)
with $\varphi^{\prime}$ given by Eq. 25.
@@
$\alpha^{\prime}_{\mathrm{L}}{}$ $=\frac{\theta^{\prime}}{2}-\varphi^{\prime}\,$, (29)
$\alpha^{\prime}_{\mathrm{R}}{}$ $=\frac{\pi}{2}-\alpha^{\prime}_{\mathrm{L}}{}\,$, (30)
$\alpha_{\mathrm{L}}{}$ $=\frac{\theta}{2}+\varphi\,$, (31)
$\alpha_{\mathrm{R}}{}$ $=\frac{\pi}{2}-\alpha_{\mathrm{L}}{}\,.$ (32)
@@
4.2.2 Connecting elementary cells in two dimensions